- Press Release
- August 19, 2022
Space Station On-Orbit Status 21 May 2003
All ISS systems continue to function nominally, except as noted previously or below.
Science Officer Ed Lu set up and then took the periodic O-OHA (on-orbit hearing assessment) test, a NASA EHS (environmental health systems) examination to assess the efficacy of acoustic countermeasures. CDR Yuri Malenchenko followed suit. [The O-OHA test involves minimum audibility measurements for each ear over a wide range of frequencies and sound pressure levels, with the crewmembers using individual-specific Prophonics earphones, Bose ANC headsets and the SLM (sound level meter). To conduct the testing, the experimenter is supported by special “EarQ” software on the MEC (medical equipment computer). The baseline test is required for about Flight Day 14 for each new Expedition and is then performed once per month.]
In line with today’s focus on acoustics, Ed Lu unstowed the joint US/Russian sound level meter (SLM) hardware and during the day took readings at 41 different locations in the Lab, Node, Airlock, FGB, Service Module (SM) and Docking Compartment (DC-1), following instructions uplinked from TsUP/MCC-M for the current Week 3.
Yuri Malenchenko continued the latest round of periodic preventive maintenance of Russian segment (RS) ventilation systems, starting out with cleaning the Group A fans and grilles in the SM (last time done: 3/18), then working in the “Pirs” DC-1 to change out its two dust filters (PF1 & PF2) and clean the mesh screens of the V1 & V2 ventilator fans (last time done: 4/21).
Later, Malenchenko performed the periodic functional closure test of the Vozdukh CO2 removal system’s emergency vacuum valves (AVK, last time done: 4/24). [The AVKs are critical because they close the Vozdukh’s vacuum access lines in the event of a malfunction in the regular vacuum valves (BVK) or a depressurization in the Vozdukh valve panel (BOA). Access to vacuum is required to vent carbon dioxide during the regeneration of the absorbent cartridges (PP). During nominal operation, the AVK valves remain open.]
Yuri also terminated the 24-hr. regeneration cycle for adsorbent bed #2 of the BMP harmful impurities filtration unit. Both filter channels are now back in Purify mode.
FE/SO Lu activated the MSG (Microgravity Science Glovebox) for its next session of the InSPACE (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions) experiment, on CA005 (coil assembly #5) with 0.4 micron particles in suspension. The MSG was later powered down again. [The test run was performed at the highest field strength for this CA (1.2 amps), and the applied magnetic field strength of 1.048A was reduced during the second monitor activity, with frequency remaining at 10 Hz. Ed monitored the runs, switching modes, sweeping the camera through the field of view and focusing on the image.]
Malenchenko started another cycle on the BIO-5 Rasteniya-2 (“Plants-2”) zero-G plant growth experiment, planting new pea seeds in the root tray of the Lada-2 greenhouse and configuring the payload for cultivation. [Rasteniya-2, a project of the Institute of Biomedical Problems (IBMP) in Moscow, is using peas of two types, a flagellate variety with reds flowers (up to 27 cm high), and an acacia-leaf variety with white flowers (up to 20 cm high). Both are dwarf peas from the plant collection of the Moscow University Genetics Department, and both have meanwhile developed pods with ripening seeds. Malenchenko’s task is to re-plant the space seeds in the Lada to grow a second generation of space peas for genetic research.]
The crew conducted an inventory audit of Russian air sampling tools available on board, checking the equipment off against an uplinked list of 33 different items for the AK-1M and AK-1MF.kits, as well as for the cartridge belts with IPD Draeger tube air samplers.
Ed Lu terminated the recharge process for the second EMU/spacesuit battery, initiated yesterday for battery maintenance.
The CDR completed the regular periodic inspection of the BRPK-1 air/liquid condensate separator in the SM.
Yuri also unstowed and set up the equipment for tomorrow’s periodic Russian MO-10 “Hematokrit” (red blood cell count) testing.
Ed performed the daily routine maintenance of SOZh life support systems, while Yuri prepared the daily IMS (inventory management system) “delta” file for updating the IMS database.
At 10:15am EDT, the crew conducted two PAO events. First, they participated in a pair of media interviews, one with the BBC “World Service” Radio, the other with WHEC-TV in Rochester, NY. The second event was a downlink message for Memorial Day, honoring the STS-107 Columbia crew.
For Russian PAO, the crew was also scheduled to downlink two TV messages for taping at MCC-Moscow and later replay. The first was an address to the residents of the ancient Russian city of Yaroslavl for their day of celebration on 5/23 (Yaroslavl, one of the most beautiful cities of Russia’s “Golden Ring”, is nearly one-thousand years old). The second message was addressed to Veteran Cosmonauts Valentina Vladimirovna Tereshkova and Valery Fedorovich Bykovsky on the occasion of the upcoming 40th anniversary (on 6/19) of their concurrent (twin) flights of two spaceships, Vostok 5 and Vostok 6.
For the on-board emergency book (EMER-1), the crew has now printed out 40 change pages and updated the old version of EMER-1, thereby recovering the third copy that was inadvertently sent down on Soyuz 5S. They now have the full complement of three emergency books required for ISS operations (one each in USOS, RS, and Soyuz TMA-2).
The PCU-2 (Plasma Contactor Unit-2) firmware controller health flag which annunciated over the weekend, has now been reset via ground command and the message cleared. PCU-2 continued to operate in Discharge mode with no other problems.
ISS continues in LVLH -YVV attitude, with U.S. CMGs in control, using less than 5% of their total momentum capability. The transition maneuver from XPOP on 5/17 required approximately 3 kg of RS thruster props.
Today’s CEO (crew earth observations) targets, now no longer restricted by the Lab science window ruled off-limit due to flight attitude, were Amazon Basin (Dynamic event. Particularly clear conditions. Recommended: Oblique mapping swath looking toward the sunglint disc [left of track] — to document detail of changeable river positions and river patterns in the major basins on the south side of the Amazon), Cape Town, South Africa (looking a touch left of track for regional views of this major port. The city has grown from its center on Table Bay [at the foot of 3000’ Table Mountain] eastwards onto the “Cape Flats” where most people now live in vast sprawling shantytowns [around and east of the international airport which may have been a visual cue for the crew]. The crew may have been able to get the city in one or two images only), Patagonian Glaciers (unusual clearing of cloud on the Pacific side of the southern Andes. Shooting glacier tongues on that side. [The same phenomenon of a local high pressure system setting up over the southern Andes occurred in January, yielding excellent photos.]), Salt Lake City, Utah (night target; nadir pass), Shanghai, China (night target; nadir pass), Riyadh, Saudi Arabia (night target; nadir pass), Alexandria, Egypt (night target; nadir pass), and Lisbon, Portugal (night target; nadir pass).
CEO images can be viewed at the websites
U.S. and Russian Segment Status (as of 12:30pm EST).
Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
- Elektron O2 generator is powered On (16 Amp mode). Vozdukh CO2 scrubber is On (manual mode 3). U.S. CDRA CO2 scrubber is Off. TCCS (trace contaminant control subsystem) is operating. MCA (major constituents analyzer) is operating. BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On; SKV-2 is Off.
- SM Working Compartment: Pressure (mmHg) — 742; temperature (deg C) — 27.2; ppO2 (mmHg) — data invalid; ppCO2 (mmHg) — data invalid.
- SM Transfer Compartment: Pressure (mmHg) — 751; temperature (deg C) — 22.0.
- FGB Cabin: Pressure (mmHg) — 748; temperature (deg C) — 21.0.
- Node: Pressure (mmHg) — 743.89; temperature (deg C) — 22.8 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- U.S. Lab: Pressure (mmHg) — 745.73; temperature (deg C) — 22.5; ppO2 (mmHg) — 172.9; ppCO2 (mmHg) — 4.7.
- Joint Airlock (Equip. Lock): Pressure (mmHg) — 745.73; temperature (deg C) — 22.5; shell heater temp (deg C) — 24.8, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- PMA-1: Shell heater temp (deg C) — 21.8.
- PMA-2: Shell heater temp (deg C) — 13.3.
(n/a = data not available)
Propulsion System (PS):
- Total propellant load available: 3576 kg (7884 lb) as of 5/15 [SM(774) + FGB(2550) + Progress(252)]. (Capability: SM — 860 kg; FGB — 6120 kg).
Electrical Power Systems (EPS):
- Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both “directed” (non-solar tracking) mode (“sun slicer” position, for drag reduction).
- SM batteries: Battery #8 is failed; batteries #1 and #4 are in ROM (capacity restoration mode); all other batteries (5) are in “Partial Charge” mode (battery #2 is exhibiting degraded performance).
- FGB batteries: Battery #4 is disconnected; all other batteries (5) are in “Partial Charge” mode.
- Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Discharge mode.
Command & Data Handling Systems:
- C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
- GNC-1 MDM is prime; GNC-2 is Backup.
- INT-1 is operating; INT-2 is Off.
- EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
- LA-1, LA-2 and LA-3 MDMs are all operating.
- PL-2 MDM is On (primary); PL-1 MDM is On
- APS-1 (automated payload switch #1) and APS-2 are both On.
- SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational. Being upgraded.
- SM Central Computer (TsVM): 3 redundant lanes (of 3) operational. Being upgraded.
Attitude Control Systems:
- 3 CMGs on-line (CMG-1 failed).
- State vector source — U.S. SIGI-1 (GPS)
- Attitude source — U.S. SIGI-2 (GPS)
- Angular rate source — RGA-2
- LVLH -YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, -y-axis in velocity vector [yaw: -90 deg, pitch: -6.8 deg, roll: -3.4 deg]), with CMG/TA (thruster assist) Momentum Management.
- Solar Beta angle: 71.0 deg (magnitude increasing).
Communications & Tracking Systems:
- FGB MDM-1 is powered Off; FGB MDM-2 is operational.
- All other Russian communications & tracking systems are nominal.
- S-band is operating nominally (on string 2).
- Ku-band is operating nominally.
- Audio subsystem is operating nominally [IAC-1 (internal audio controller #1) being analyzed after self-test error. IAC-2 is prime.]
- Video subsystem operating nominally.
- HCOR (high-rate communications outage recorder) is operating nominally.
- SSRMS/Canadarm2 based at MBS PDGF #1 with Keep Alive (KA) power on both strings.
- MBS: KA power on both strings.
- MT: latched at WS4, with KA power.
- POA: KA power on both strings.
- RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.
ISS Orbit (as of this morning, 7:01am EDT [= epoch]):
- Mean altitude — 389.4 km
- Apogee — 393.4 km
- Perigee — 385.4 km
- Period — 92.34 min.
- Inclination (to Equator) — 51.63 deg
- Eccentricity — 0.0005927
- Orbits per 24-hr. day — 15.59
- Solar Beta Angle — 71.0 deg (magnitude increasing)
- Mean altitude loss in last 24 hours — 75 m
- Revolutions since FGB/Zarya launch (Nov. ’98) — 25683
- For more on ISS orbit and worldwide naked-eye visibility dates/times, see